### v1

@@ 1 Introduction
The ability to seamlessly coordinate whole-body movements while navigating complex environments represents one of humanity’s most remarkable capabilities [1, 2]. From squatting to retrieve objects from the ground to walking across rooms while carrying items, humans effortlessly integrate locomotion and manipulation in ways that remain challenging for robots [3, 4, 5, 6, 7]. Humanoid robots (humanoids henceforth), with their human-like morphology, offer the promise of replicating these capabilities—potentially enabling applications from household assistance to operations in hazardous environments where human-like dexterity and mobility are essential [8, 9, 10, 11, 12].
@@ 2 Related Work
Long-horizon task execution [49] has been extensively studied for fixed-base arms [50, 51, 52, 53], mobile manipulators [54, 55, 56, 57, 58], and aerial manipulators [59], typically in structured settings. In contrast, humanoid teleoperation remains limited to short-horizon motion replication [19, 23, 24], operating open-loop due to difficulties in real-time global state estimation for bipedal systems. Although recent advances in odometry have improved state tracking for legged robots [60, 61], their application to long-horizon humanoid control remains largely unexplored. To bridge this gap, we integrate LiDAR odometry into our teleoperation framework to enable closed-loop error correction and significantly reduce accumulated drift.
@@
Large-scale motion capture (MoCap) datasets [29, 62] have been instrumental in training humanoid control policies [15, 19, 23, 63]. Even after augmenting the datasets with generative models [24], these datasets were still confined primarily to animation and graphics [64, 25] rather than robotics applications. While they contain semantically distinct actions (e.g., waving, hugging, drinking), they underrepresent the kinematic configurations and dynamic transitions required for robust, generalizable controller training in real-world scenarios. To address these limitations, we introduce Cloned by augmenting AMASS [29] through motion editing and collecting additional human MoCap data, specifically tailored for humanoid controllers. This expansion increases coverage of motions and transitions relevant to humanoid control tasks.
@@ 3 The Clone Framework
Our teleoperation framework, Clone, captures a minimal set of control signals from the teleoperator, consisting solely of the 6D poses (position and orientation) of both wrists and the 3D position of the head, tracked using an AVP headset. These three points (see also Fig. 3) serve as the complete control interface, providing an intuitive yet powerful means of directing the humanoid’s whole-body motion while maintaining a simple setup that requires no additional hardware or complex calibration procedures.
@@ 3.4 The Cloned Dataset
The training dataset Cloned comprises three complementary components to support robust whole-body teleoperation. These include: (i) an augmented AMASS [29] subset of 149 curated sequences featuring diverse pairings of upper- and lower-body movements, enhanced via targeted motion editing to increase compositional diversity and policy generalization; (ii) 14 custom sequences captured with an IMU-based Xsens MoCap system to fill coverage gaps, emphasizing continuous transitions and diverse upper-body poses critical for manipulation; and (iii) systematic hand orientation augmentation through procedurally generated 6D wrist targets, smoothed via Spherical Linear Interpolation (SLERP) to ensure coherent and natural hand motions for teleoperation.
@@ 4 Real-World Experiments
To evaluate global positioning accuracy over extended distances, we designed a controlled path-following experiment. We marked fixed initial positions for both operator and robot, then had the operator walk along straight paths to target positions at $3$m, $6$m, and $8.9$m distances while teleoperating the robot. We measured the discrepancy between the robot’s final position and expected position as tracking error, repeating each condition ten times.
Clone achieved a mean tracking error of $5.1$cm with a maximum deviation of $12.0$cm at $8.9$m, as shown in Fig. 4. This exceptional positional accuracy validates that Clone’s closed-loop error correction effectively mitigates accumulated positional errors during extended teleoperation.
Statistical analysis confirms consistent tracking performance across all tested distances. Independent samples t-tests comparing distance groups yielded: $3$m versus $6$m ($t=0.165,p=0.871$) and $6$m versus $8.9$m ($t=0.048,p=0.963$). With all p-values $>0.05$, Clone shows no significant performance degradation across the tested range, demonstrating that our approach maintains high accuracy over extended distances without drift accumulation.
@@
To validate system performance in complex scenarios, we conducted extended teleoperation sessions incorporating multiple movement types. As visualized in Fig. 6, we recorded a continuous teleoperation sequence where the operator traversed a complex path spanning over $15$m, incorporating diverse locomotion patterns including forward walking, turning, side-stepping, and returning to the original position.
@@ 5 Simulations
Method $\mathbf{SR}$uparrow $E_{\text{mpkpe}}$downarrow $E_{\text{r-mpkpe}}$downarrow $E_{\text{vel}}$downarrow $E_{\text{hand-rot}}$downarrow
Clone $\dagger$ $100\%$ $113.97$ $35.55$ $245.11$ $4.73$
@@
To assess Clone’s robustness across varying postures, we evaluated performance in tracking motions with head heights from $1.2$m (standing) to $0.6$m (deep squatting) in $0.1$m decrements. We generated these challenging motions by systematically editing sequences from the Cloned dataset, creating unseen poses that test teleoperation system limits.
@@
6 Conclusion and Discussions
We present Clone, a closed-loop MoE-based teleoperation system that enables comprehensive humanoid control while addressing accumulated tracking errors in long-horizon tasks. Our approach integrates three key innovations: an MoE architecture coordinating diverse motion patterns, LiDAR-based closed-loop error correction preventing positional drift, and the Cloned dataset providing robust whole-body coordination training. Experimental validation demonstrates exceptional performance: precise global position tracking with 5.1cm mean error over 8.9m distances, accurate whole-body motion coordination across diverse skills, and robust long-horizon teleoperation using only minimal MR headset input.
While demonstrating significant capabilities, important limitations require future investigation. The minimal input configuration, though enabling intuitive control with head and hand tracking, inherently limits fine-grained stability control in certain scenarios. Additionally, the system exhibits reduced performance during highly dynamic movements like jumping, stemming from training data constraints and balance control challenges.
@@ A.1 Formulation
We formulate the humanoid teloperation as a Markov Decision Process (MDP) =$\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\}$. $\mathcal{S}$ includes proprioceptive states $s$ and task-oriented observations $o^{task}$. The action space $\mathcal{A}\in\mathcal{R}^{29}$ represents the humanoid’s joint angles in our method. $\mathcal{T}$ is the transition function conditioned on $\mathcal{S}$ and $\mathcal{A}$. The reward functions $\mathcal{R}$ is defined based on $\mathcal{S},\mathcal{A}$. A policy $\pi$ is proposed to maximize the overall reward $\mathcal{R}$ using the Proximal Policy Optimization (PPO) algorithm.
@@ Appendix B Reward Functions and Domain Randomization
Torque $\lVert\boldsymbol{\tau}\rVert_{2}^{2}$ $-0.0001$
Torque limits $[{\tau\notin[{\tau}_{\text{min}},{\tau}_{\text{max}}]}]_{1}$2
DoF position limits $[{\mathbf{p}_{t}\notin[\mathbf{p}_{\text{min}},\mathbf{p}_{\text{max}}]}]_{1}$625
DoF velocity limits $[{\mathbf{\dot{p}}_{t}\notin[\mathbf{\dot{p}}_{\text{min}},\mathbf{\dot{p}}_{\text{max}}]}]_{1}$50
Termination $\text{termination}_{1}$ $-e^{4}$
DoF acceleration $\lVert\mathbf{\ddot{q}_{t}}\rVert_{2}^{2}$ $-1.1e^{-5}$
DoF velocity $\lVert\mathbf{\dot{q}_{t}}\rVert_{2}^{2}$ $-0.004$
Lower-body action rate $\lVert\mathbf{a}_{t}^{\text{lower}}-\mathbf{a}_{t-1}^{\text{lower}}\rVert_{2}^{2}$ $-1.0$
Upper-body action rate $\lVert\mathbf{a}_{t}^{\text{upper}}-\mathbf{a}_{t-1}^{\text{upper}}\rVert_{2}^{2}$ $-0.3$
@@
Stumble $[(\mathbf{F}_{\text{feet}}^{xy}>5\times\mathbf{F}_{\text{feet}}^{z})]_{1}$ $-1.25e^{4}$
Slippage $\lVert v^{\text{feet}}_{t}\rVert_{2}^{2}\cdot[(\mathbf{F}_{\text{feet}}\geq 1)]_{1}$ $-80$
Feet orientation $\lVert\mathbf{R}_{z}^{\text{feet}}\rVert$ $-62.5$
In the air $[(\mathbf{F}_{\text{feet}}^{\text{left}},\mathbf{F}_{\text{feet}}^{\text{right}}<1)]_{1}$ $-750$
Orientation $\lVert\mathbf{R}_{z}^{\text{root}}\rVert$ $-50$
@@
Body rotation $\exp(-0.1\lVert\mathbf{\theta}\ominus\mathbf{\hat{\theta}}\rVert)$ $400$
@@
Body angular velocity $\exp(-0.01\lVert\boldsymbol{\omega}-\boldsymbol{\hat{\omega}}\rVert_{2})$ $8$
@@
[…]
@@ Appendix C Implement Details
We train our policy in IsaacGym using a single A800800800800 GPU. The teacher policy is trained for $1M$ iterations with $8192$ parallel environments, while the student policy is trained for $600K$ iterations with $4096$ parallel environments. In our implementation, the student policy consists of $L=3$ MoE layers, each containing $N=4$ experts. It uses a history length of $H=25$ frames and activates the top $k=2$ experts based on the highest weights determined by the router.
@@ D.1 Evaluation Metrics
We evaluated Clone on motion tracking tasks from Cloned using five metrics: success rate $\mathbf{SR}$ (%), mean per-keybody position error (MPKPE) $E_{\text{mpkpe}}$ (mm), root-relative mean per-keybody position error (R-MPKPE) $E_{\text{r-mpkpe}}$ (mm), average joint velocity error $E_{\text{vel}}$ (mm/s), and hand orientation tracking error $E_{\text{hand}}$. Success rate ($\mathbf{SR}$) represents the proportion of episodes where: (i) the robot maintains balance without falling, and (ii) the average per-keybody distance between the robot and reference motion remains below $1.5$m across the three controlled joints. We defined the hand orientation tracking error as $E_{\text{hand}}=1-\left<\hat{\textrm{q}},\textrm{q}\right>^{2}$, where $\hat{\textrm{q}}$ and q represent the reference and robot hand quaternions.

### v2

@@ 1 Introduction
The ability to seamlessly coordinate whole-body movements while navigating complex environments represents one of humanity’s most remarkable capabilities [1, 2]. From squatting to retrieve objects from the ground to walking across rooms while carrying items, humans effortlessly integrate locomotion and manipulation in ways that remain challenging for robots [3, 4, 5, 6, 7]. Humanoid robots (humanoids henceforth), with their human-like morphology, offer the promise of replicating these capabilities—potentially enabling applications from household assistance to operations in hazardous environments where human-like dexterity and mobility are essential [8, 9, 10, 11].
@@ 2 Related Work
Long-horizon task execution [48] has been extensively studied for fixed-base arms [49, 50, 51, 52], mobile manipulators [53, 54, 55, 56, 57], and aerial manipulators [58], typically in structured settings. In contrast, humanoid teleoperation remains limited to short-horizon motion replication [18, 22, 23], operating open-loop due to difficulties in real-time global state estimation for bipedal systems. Although recent advances in odometry have improved state tracking for legged robots [59, 60, 61], their application to long-horizon humanoid control remains largely unexplored. To bridge this gap, we integrate LiDAR odometry into our teleoperation framework to enable closed-loop error correction and significantly reduce accumulated drift.
@@
Large-scale motion capture (MoCap) datasets [28, 62] have been instrumental in training humanoid control policies [14, 18, 22, 63, 64]. Even after augmenting the datasets with generative models [23], these datasets were still confined primarily to animation and graphics [65, 24] rather than robotics applications. While they contain semantically distinct actions (e.g., waving, hugging, drinking), they underrepresent the kinematic configurations and dynamic transitions required for robust, generalizable controller training in real-world scenarios. To address these limitations, we introduce Cloned by augmenting AMASS [28] through motion editing and collecting additional human MoCap data, specifically tailored for humanoid controllers. This expansion increases coverage of motions and transitions relevant to humanoid control tasks.
@@ 3 The Clone Framework
Our teleoperation framework captures a minimal set of control signals from the teleoperator, consisting solely of the 66D poses (position and orientation) of both wrists and the 33D position of the head, tracked using an AVP headset. These three points (see also Fig. 3) serve as the complete control interface, providing an intuitive yet powerful means of directing the humanoid’s whole-body motion while maintaining a simple setup that requires no additional hardware or complex calibration.
@@ 3.4 The Cloned Dataset
The training dataset Cloned comprises three complementary components to support robust whole-body teleoperation. These include: (i) an augmented AMASS [28] subset of 149 curated sequences featuring diverse pairings of upper- and lower-body movements, enhanced via motion editing to increase compositional diversity and policy generalization(see Sec. C.1); (ii) 14 custom sequences captured with an IMU-based Xsens MoCap system to fill coverage gaps, emphasizing continuous transitions and diverse upper-body poses critical for manipulation; and (iii) systematic hand orientation augmentation through procedurally generated 6D wrist targets, smoothed via Spherical Linear Interpolation (SLERP) to ensure coherent and natural hand motions for teleoperation.
@@ 4 Real-World Experiments
To evaluate global positioning accuracy over extended distances, we designed a controlled path-following experiment. (i) Straight-path tracking. We fixed the initial positions of both the operator and the robot. The operator then walked along straight paths to target positions at 33m, 66m, and 8.98.9m while teleoperating the robot. We measured the discrepancy between the robot’s final and expected positions as the tracking error and repeated each condition ten times. (ii) Curved-path tracking. We teleoperated the humanoid along a 1010m trajectory with two $90^{\circ}$ turns, representative of typical household paths, and repeated the trial six times to measure translational and rotational tracking errors.
Clone achieved a mean tracking error of $5.1$ cm in straight-path tracking, with a maximum deviation of $12.0$ cm at 8.98.9m (see Fig. 4). These results show that Clone’s closed-loop error correction effectively mitigates accumulated errors during extended teleoperation.
Statistical analysis confirms consistent tracking performance across all tested distances. Independent samples t-tests comparing distance groups yielded: 33m versus 66m ($t=0.165,p=0.871$) and 66m versus 8.98.9m ($t=0.048,p=0.963$). With all p-values >0.05>0.05, Clone shows no significant performance degradation across the tested range, demonstrating that our approach maintains high accuracy over extended distances without drift accumulation.
In curved-path tracking, the mean error is $20$ cm (maximum $27$ cm), and the mean rotational drift was $2^{\circ}$ between the operator’s and humanoid’s orientations. These results highlight Clone ’s ability to sustain high precision even over long and complex trajectories.
@@
To validate system performance in complex scenarios, we conducted extended teleoperation sessions incorporating multiple movement types. As visualized in Fig. 6, we recorded a continuous teleoperation sequence where the operator traversed a complex path spanning over 1515m, incorporating diverse locomotion patterns including forward walking, turning, side-stepping, and returning to the original position.
@@ 5 Simulations
Method $\mathbf{SR}$ $\uparrow$ $E_{\text{mpkpe}}$ $\downarrow$ $E_{\text{r-mpkpe}}$ $\downarrow$ $E_{\text{vel}}$ $\downarrow$ $E_{\text{hand-rot}}$ $\downarrow$
Clone $\dagger$ $00\% 11$3.97113.97 $35.55$ $245.11$ $4.73$
@@
To assess Clone’s robustness across varying postures, we evaluated performance in tracking motions with head heights from 1.21.2m (standing) to 0.60.6m (deep squatting) in 0.10.1m decrements. We generated these challenging motions by systematically editing sequences from the Cloned dataset, creating unseen poses that test teleoperation system limits.
@@
6 Conclusion
We present Clone, a closed-loop MoE-based teleoperation system that enables comprehensive humanoid control while addressing accumulated tracking errors in long-horizon tasks. Our approach integrates three key innovations: (i) an MoE architecture that coordinates diverse motion skills while maintaining natural upper- and lower-body coordination, (ii) LiDAR-based closed-loop error correction that prevents positional drift accumulation through real-time feedback, (iii) and the Cloned dataset that augments AMASS with hand orientations and additional motion-captured sequences for robust whole-body coordination training. Experimental validation demonstrates exceptional performance: $5.1$ cm mean global position tracking error over $8.9$ m trajectories, accurate whole-body coordination across diverse skills including complex coordinated movements like object retrieval from ground level, and robust long-horizon teleoperation requiring only head and hand tracking from a single commercial MR headset.
7 Limitations
While Clone demonstrates significant capabilities, it still has important limitations that warrant further investigation. The minimal input configuration, though enabling intuitive control with head and hand tracking, inherently limits fine-grained stability control in certain scenarios. Additionally, the system exhibits reduced performance during highly dynamic movements like jumping, stemming from training data constraints and balance control challenges.
@@ A.1 Formulation
We formulate the humanoid teloperation as a Markov Decision Process (MDP) $\mathcal{M}=\{\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\}$. $\mathcal{S}$ includes proprioceptive states $s$ and task-oriented observations $o^{task}$. The action space $\mathcal{A}\in\mathcal{R}^{29}$ represents the humanoid’s joint angles in our method. $\mathcal{T}$ is the transition function conditioned on $\mathcal{S}$ and $\mathcal{A}$. The reward functions $\mathcal{R}$ is defined based on $\mathcal{S},\mathcal{A}$. A policy $\pi$ is proposed to maximize the overall reward $\mathcal{R}$ using the Proximal Policy Optimization (PPO) algorithm.
@@ Appendix B Reward Functions and Domain Randomization
Torque $\lVert\boldsymbol{\tau}\rVert_{2}^{2}$ 0.0001-0.0001
Torque limits ${\tau\notin[{\tau}_{\text{min}},{\tau}_{\text{max}}]}]_{1} -$2
DoF position limits ${\mathbf{p}_{t}\notin[\mathbf{p}_{\text{min}},\mathbf{p}_{\text{max}}]}]_{1} -$625
DoF velocity limits ${\mathbf{\dot{p}}_{t}\notin[\mathbf{\dot{p}}_{\text{min}},\mathbf{\dot{p}}_{\text{max}}]}]_{1} -$50
Termination $\text{termination}_{1}$ $e^{4}$
DoF acceleration $\lVert\mathbf{\ddot{q}_{t}}\rVert_{2}^{2}$ $1.1e^{-5}$
DoF velocity $\lVert\mathbf{\dot{q}_{t}}\rVert_{2}^{2}$ 0.004-0.004
Lower-body action rate $\lVert\mathbf{a}_{t}^{\text{lower}}-\mathbf{a}_{t-1}^{\text{lower}}\rVert_{2}^{2}$ 1.0-1.0
Upper-body action rate $\lVert\mathbf{a}_{t}^{\text{upper}}-\mathbf{a}_{t-1}^{\text{upper}}\rVert_{2}^{2}$ 0.3-0.3
@@
Stumble $(\mathbf{F}_{\text{feet}}^{xy}>5\times\mathbf{F}_{\text{feet}}^{z})]_{1}$ $1.25e^{4}$
Slippage $\lVert v^{\text{feet}}_{t}\rVert_{2}^{2}\cdot[(\mathbf{F}_{\text{feet}}\geq 1)]_{1}$ 80-80
Feet orientation $\lVert\mathbf{R}_{z}^{\text{feet}}\rVert$ 62.5-62.5
In the air $(\mathbf{F}_{\text{feet}}^{\text{left}},\mathbf{F}_{\text{feet}}^{\text{right}}<1)]_{1}$ 750-750
Orientation $\lVert\mathbf{R}_{z}^{\text{root}}\rVert$ 50-50
@@
[…]
@@
Body angular velocity $\exp(-0.01\lVert\boldsymbol{\omega}-\boldsymbol{\hat{\omega}}\rVert_{2})$ 88
@@
Torque RFI $0.05\times\text{torque limit}\ \text{N}\cdot\text{m}$
@@ Appendix C Implement Details
C.1 Data Augmentation
In our implementation, we filter out physically infeasible AMASS data and select motions with large upper- and lower-body workspaces as oracle motions. These motions are further modified by concatenating body parts or accelerating sequences.
C.2 Model Architecture
The student policy is composed of $L=3$ MoE layers, each containing $N=4$ experts, where each expert is implemented as an MLP with dimensions ($2048$, $512$, $512$, $256$). The policy uses a history length of $H=25$ frames and activates the top $k=2$ experts based on the highest weights determined by the router. The AMP discriminator is a $3$-layer MLP ($256$, $256$, $256$) that is updated online during training on Cloned. For comparison, the baseline model Clone $\dagger$ uses a single MLP with architecture ($2048$, $1024$, $512$, $512$).
C.3 Policy Training
We train our policy in IsaacGym using a single A800800 GPU. The teacher policy is trained for 11M iterations with $8192$ parallel environments, while the student policy is trained for $600K$ iterations with $4096$ parallel environments. Training the teacher policy requires $\sim 480$K simulation steps ($\sim 20$K PPO steps) and $\sim 24$ hours on a single A800800 GPU. The student policy requires $\sim 48$ hours on a single $3090$ Ti.
@@ D.1 Evaluation Metrics
We evaluated Clone on motion tracking tasks from Cloned using five metrics: success rate $\mathbf{SR}$ (%), mean per-keybody position error (MPKPE) $E_{\text{mpkpe}}$ (mm), root-relative mean per-keybody position error (R-MPKPE) $E_{\text{r-mpkpe}}$ (mm), average joint velocity error $E_{\text{vel}}$ (mm/s), and hand orientation tracking error $E_{\text{hand}}$. Success rate ($\mathbf{SR}$) represents the proportion of episodes where: (i) the robot maintains balance without falling, and (ii) the average per-keybody distance between the robot and reference motion remains below 1.51.5m across the three controlled joints. We defined the hand orientation tracking error as $E_{\text{hand}}=1-\left<\hat{\textrm{q}},\textrm{q}\right>^{2}$, where $\hat{\textrm{q}}$ and q represent the reference and robot hand quaternions.